What is Ecell for the following hypothetical reaction exactly as written (in volts) at T = 25°C given the conditions below? Express your answer to 2 decimal places.

X(s) + 2 Y+(aq) → X2+(aq) + 2 Y(s)

cell = -1.38 V

[Y+] = 0.001 mol L-1

[X2+] = 1.33 mol L-1
Answer #1


Number of electron being transferred in balanced reaction is 2
So, n = 2

use:
E = Eo - (2.303*RT/nF) log {[X2+]^1/[Y+]^2}

Here:
2.303*R*T/F
= 2.303*8.314*298.0/96500
= 0.0591

So, above expression becomes:
E = Eo - (0.0591/n) log {[X2+]^1/[Y+]^2}
E = -1.38 - (0.0591/2) log (1.33^1/0.001^2)
E = -1.38-(0.181)
E = -1.561 V
Answer: -1.56 V